Initial 100 milligram of a radio active substance was present.After 6 hour the mass has been decreased by 3%. If the rate of decay is proportional to the amount of the substance present at time t,find the amount remaining after 24hours.


1
Expert's answer
2021-11-30T07:24:50-0500

Let "A(t)=" the amount of radio active substance that re,aims after "t" hours in

milligram.

The rate of decay is proportional to the amount of the substance present at time "t"


"\dfrac{dA}{dt}=kA"

"\dfrac{dA}{A}=ktdt"

Integrate


"\int \dfrac{dA}{A}=\int ktdt"

"A(t)=Ce^{kt}"

Initial 100 milligram of a radio active substance was present


"A(0)=C=100"

Then


"A(t)=100e^{kt}"

After 6 hour the mass has been decreased by 3%


"e^{6t}=1-0.3"

"t=\dfrac{\ln 0.97}{6}"

Then


"A(t)=100e^{({\ln 0.97 \over 6})t}"

"A(t)=100(0.97)^{t\/6}"

"A(24)=100(0.97)^{24\/6}=100(0.97)^4"

"=88.529281\approx88.53 \ (mg)"

88.53 milligram of a radio active substance remained after 24 hours.